OpenStudy (anonymous):

when 15 is divided by a number,the result is 0.3 find the number

OpenStudy (mathhelp346):

4.5

OpenStudy (smurfy14):

answer is 50

OpenStudy (anonymous):

15/x=3/10 and then cross multiply

OpenStudy (anonymous):

50 mate.

OpenStudy (anonymous):

150=3x x=50 smurfy is correct

OpenStudy (smurfy14):

15/n=0.3 which means 15=0.3n so divide 15 by 0.3 and you get 50.

jhonyy9 (jhonyy9):

50 is right